20cm3 of gaseous hydrocarbon z was burnt in 210cm3 of excess oxygen, after cooling to room temperature the gaseous product occupied a volume of 170cm3. When the gaseous mixture was passed through conc potassium hydroxide solution the volume reduced by 80cm3.

1. Write the equation for the reaction that took place

2.Determine molecular formular of z.


1
Expert's answer
2020-11-02T08:50:57-0500

1) equation: let the hydrocarbon be CxHy

CxHy + y/4 O2 = ×CO2 + y/2H2O


2)moles of CO2 produced = 1×80/24000=0.00333 moles

Moles of H2O produced = 1×90/24000= 0.00375

X/y/2 = 0.00333/0.00375

Y= 8

The molecular formula is therefore C4H8